OpenStudy (anonymous):

A soccer ball is kicked toward the goal. The height of the ball is modeled by the function h(t) = -16t2 + 48t where t equals the time in seconds and h(t) represents the height of the ball at time t seconds. What is the axis of symmetry, and what does it represent?

OpenStudy (anonymous):

can I get a straight up answer?

OpenStudy (anonymous):

t = 1.5; it takes 1.5 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 1.5; it takes 1.5 seconds to reach the maximum height and 1.5 seconds to fall back to the ground t = 3; it takes 3 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 3; it takes 3 seconds to reach the maximum height and 6 seconds to fall back to the ground

OpenStudy (anonymous):

you are being asked for the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{48}{2\times (-16)}=1.5\]

OpenStudy (anonymous):

i guess it is going "straight up" as you might say

OpenStudy (anonymous):

no by straight up I mean I want an answer right cause i have to finish the assignment now.

OpenStudy (anonymous):

right now I mean

OpenStudy (anonymous):

but thanks